OpenStudy (anonymous):

plz anyone help me find the limit of this [1/(8+x) -1/8]/X the limit as x goes to 0

OpenStudy (anonymous):

\[-\frac{1}{64} \]\[\frac{-\frac{1}{8}+\frac{1}{8+x}}{x}=-\frac{1}{8 (8+x)}=-\frac{1}{(64+8 x)} \]